A particular bulb has lifetimes of 1500 hours with a standard deviation of 6 hours. Find a range in which it can be guaranteed that 93.75% of the lifetimes of this brand of car lie.



(1489 and 1512)



(1482 and 1518)



(1476 and 1524)



(1470 and 1530)

1
Expert's answer
2021-11-12T12:29:44-0500

Mean = 1500

Standard deviation = 6

Z value at 93.75% confidence interval = 1.534


"Mean\pm Z \frac{s}{\sqrt{n}}"


"1500\pm 1.534 \frac{6}{\sqrt{1}}"


"1500\pm 1.534 \times6"


(1489 and 1512)